OpenStudy (anonymous):

Find the value of y so that the line passing through (9, 0) and (3, y) has a slope of negative one half

myininaya (myininaya):

\[\frac{y-0}{3-9}=\frac{-1}{2}\]

myininaya (myininaya):

all we need to do is solve this for y

myininaya (myininaya):

\[\frac{y}{-6}=\frac{-1}{2}\]

myininaya (myininaya):

\[2y=6 \] i crossed multiplied

myininaya (myininaya):

now divide both sides by what and you will get y

OpenStudy (anonymous):

2?

myininaya (myininaya):

yep

myininaya (myininaya):

and so y=6/2=3 gj